OpenStudy (anonymous):

Hey guys I will Medal and Fan, I have a test tomorrow and I was wondering if somebody could run me through on how to find the 6 trig functions from the Unit circle?

OpenStudy (anonymous):

I know mostly but I need a review.

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

yay @Nnesha

OpenStudy (anonymous):

can you help??

Nnesha (nnesha):

2 ways) learn the pattern /memorize familiar with 45-45-90 / 30-60-90 triangle i have 10 mints then i gtg somwhere :D

Nnesha (nnesha):

btw are you allowed to use unit circle ?

OpenStudy (anonymous):

ok I know how to get sine and cosine but what about the other 4

OpenStudy (anonymous):

no but I know how to get the coordinates without it

Nnesha (nnesha):

ohhh i know what you are asking now they are reciprocal of each other let's say sin is 2/3 then csc would be 3/2

OpenStudy (anonymous):

ok cool then the only thing i need is tangent, how do i get tangent and cotangent

Nnesha (nnesha):

\[\large\rm sin \theta = \frac{ 1 }{ \csc \theta }~~~\cos \theta=\frac{ 1 }{ \sec \theta }~~~\tan \theta =\frac{ 1 }{ \cot }=\frac{ \sin }{ \cos}\]

OpenStudy (anonymous):

can you kind of quiz me, like what is the tangent of 45 degrees? just a few questions want to make sure I got it down

Nnesha (nnesha):

 tan = sin over cos

Nnesha (nnesha):

sure!

OpenStudy (anonymous):

ok hit me!

Nnesha (nnesha):

\(\huge\color{red}{\rm Q:1}\) \[\tan( 120)\] \(\huge\color{red}{\rm Q:2}\) \[\sec(270)\] \(\huge\color{red}{\rm Q:3}\) \[\csc(360)\]

Nnesha (nnesha):

degrees*

OpenStudy (anonymous):

sorry but I have to go, I get this stuff though , thanks for helping me review. you helped a lot I will medal you if you medal me.

Nnesha (nnesha):

what did you get ? :) ;o